# Topology and Spectrum in Measurement-Induced Phase Transitions

### v2

@@ Abstract
To discuss the topology of the three phases, we first investigate the Lyapunov spectrum in the circuit with OBC, finding that the critical phase corresponds to a gapless phase without Majorana zero modes. Figure 3(a) shows the $\mu_{e}$ depedence of the non-negative single-particle Lyapunov spectrum. As observed in the measurement-only case, the lowest mode $z_{1}$ become gapless (gapped) in the topological (trivial) phase. In the critical phase for $0.4\lesssim\mu_{e}\lesssim 0.6$, not only the lowest mode $z_{1}$ but also higher modes $z_{k(\neq 1)}$ appear to become gapless, implying the closing of the bulk Lyapunov gap. Interestingly, the $L$-dependence of $z_{1}$ seems different between the topological area-law phase and the critical phase; $z_{1}$ decays faster than power law in the topological area-law phase ($\mu_{e}\lesssim 0.4$), while it appears to decay in $1/L$ for our available system sizes in the critical phase ($0.4\lesssim\mu_{e}\lesssim 0.6$[53].
@@
The above observation of the bulk spectrum is also confirmed from the lowest spectrum $z_{1}$ under the PBC and APBC. Indeed, as shown in Fig. 3(b,c), $z_{1}$ goes to zero for $0.4\lesssim\mu_{e}\lesssim 0.6$, whereas the bulk gap is open for other $\mu_{e}$. While the scaling form of $z_{1}$ right at the transitions $\mu_{e}\simeq 0.4$ and $0.6$ is close to $1/L$, it appears to decay faster than $1/L$ inside the critical phase (except the APBC circuit at $\mu_{e}=0.5$, where exact gap closing is anticipated nearby) [53].
@@
The above results for the spectral gaps indicate that the bulk-edge correspondence is obscured for the gapless phase if we consider $\chi_{\infty}\equiv\lim_{T\rightarrow\infty}\chi_{T}(\bm{s})$ for finite $L$, which is confirmed to be convergent irrespective of $\bm{s}$. That is, since $z_{1}$ can only be exactly zero near $\mu_{e}=0.5$ for APBC, the asymptotic topological invariant can switch the sign only at this point for finite system sizes. Indeed, we have numerically observed that $\chi_{T}$ flows towards $-1$ for $\mu_{e}<0.5$ and $+1$ for $\mu_{e}>0.5$ as $T$ increases [53]. In contrast, as discussed above, the Majorana zero mode seems absent even for $0.50.4\lesssim\mu_{e}<$ 0.5. Therefore, while the critical phase is unstable in static Majorana chains and is thus unique to monitored Majorana chains, the bulk-edge correspondence is obscured since it corresponds to the gapless phase. This is reminiscent of the fact that topological phase is usually well defined in the gapped phase in static systems.
@@
Remarkably, however, we find that the topological invariant dynamically characterizes the different phases, including the critical gapless phase. Specifically, we consider the sample-averaged invariant, $\chi_{T}$, for the timescale $T=\mathcal{O}(L)$. Figure 3(d) shows $\chi_{T=L}$ for different system sizes, from which we find three crossings with $\mu_{e}\simeq 0.4,0.5,$ and $0.6$. The locations of the first and last crossings coincide well to the transitions between the area-law gapped phases and critical gapless phase. Notably, this characterization is attributed to the divergent relaxation time of the invariant in the gapless phase. If we denote the timescale as $\tau_{\mathrm{relax}}$, it is at least longer than $\sim 1/z_{1}$, which increases faster than $L$ inside the gapless phase. Therefore, $\chi_{T=L}$ will not converge and almost stay zero for the gapless phase. This is contrasted to the gapped phases with $z_{1}=\mathcal{O}(L^{0})$, where $\chi_{T}$ rapidly converge to ±$\pm 1$ for $T=\mathcal{O}(L)$. Note that the crossing point at $\mu_{e}\simeq 0.5$ is different from the boundaries of the three phases. The above argument indicates that this crossing is a finite-size artifact, and $\chi_{T=L}$ will flatly become zero in the entire critical phase for $L\rightarrow\infty$.
Summary and outlook.–
@@
We have shown that monitored Majorana circuits in the gapped phases exhibit the bulk-edge correspondence through investigating the Lyapunov spectrum, gapless edge modes, and topological invariants based on the fermion parity. To define the topological invariant, we have used a unique methodology where the trajectory under APBC is defined with respect to that under PBC, according to twisted measurement outcomes at the boundary. We have also shown that the critical phase corresponds to the gapless phase in terms of the Lyapunov spectrum and is dynamically characterized by the topological invariant in a timescale $t=\mathcal{O}(L)$.
@@
Our study opens the way to explore topological features in monitored systems through the Lyapunov analysis. An important future direction is to examine many-body systems, where our topological invariant based on the twisted measurement outcomes readily applies. For example, it is intriguing to analyze monitored circuits with $Z_{2}\times Z_{2}$ symmetry, for which measurements can stabilize an area-law phase with a symmetry-protected topological order of the cluster state [26, 38, 32]. Such a topological area-law phase will be charactarized by (nearly) four-fold degenerate Lyapunov spectrum under the OBC from gapless edge modes and a topological invariant measuring the change of one $Z_{2}$ charge by twisting boundary measurement outcomes for another $Z_{2}$ symmetry.
@@
Note added.— While this work was in its final stage, we learned about a recent paper which investigates a similar issue [72].
Data availability.— Data and simulation codes are available upon reasonable request.

### v4

@@ Abstract
To discuss the topology of the three phases, we first investigate the Lyapunov spectrum in the circuit with the OBC, finding that the critical phase corresponds to a gapless phase without Majorana zero modes. Figure 3(a) shows the $\mu_{o}$ dependence of the non-negative single-particle Lyapunov spectrum. As observed in the measurement-only case, the lowest mode $z_{1}$ becomes gapless (gapped) in the topological (trivial) phase. In the critical phase for $0.4\lesssim\mu_{o}\lesssim 0.6$, not only the lowest mode $z_{1}$ but also higher modes $z_{k(\neq 1)}$ appear to become gapless, implying the closing of the bulk Lyapunov gap. Interestingly, the $L$ dependence of $z_{1}$ seems different between the topological area-law phase and the critical phase; $z_{1}$ decays faster than power law in the topological area-law phase ($\mu_{o}\lesssim 0.4$), while it appears to decay in $1/L$ for our available system sizes in the critical phase ($0.4\lesssim\mu_{o}\lesssim 0.6$[55].
@@
The above observation of the bulk spectrum is also confirmed from the lowest spectrum $z_{1}$ under the PBC and APBC. Indeed, as shown in Figs. 3(b) and 3(c), $z_{1}$ goes to zero for $0.4\lesssim\mu_{o}\lesssim 0.6$, whereas the bulk gap is open for other $\mu_{o}$. While the scaling form of $z_{1}$ right at the transitions $\mu_{o}\simeq 0.4$ and $0.6$ is close to $1/L$, it appears to decay faster than $1/L$ inside the critical phase (except the APBC circuit at $\mu_{o}=0.5$, where exact gap closing is anticipated nearby) [55].
@@
The above results for the spectral gaps indicate that the bulk-edge correspondence is obscured for the gapless phase if we consider $\chi_{\infty}\equiv\lim_{T\rightarrow\infty}\chi_{T}(\bm{s})$ for finite $L$, which is confirmed to be convergent irrespective of $\bm{s}$. That is, since $z_{1}$ can only be exactly zero near $\mu_{o}=0.5$ for the APBC, the asymptotic topological invariant can switch the sign only at this point for finite system sizes. Indeed, we have numerically observed that $\chi_{T}$ flows toward $-1$ for $\mu_{o}<0.5$ and $+1$ for $\mu_{o}>0.5$ as $T$ increases [55]. In contrast, as discussed above, the Majorana zero modes seem absent even for $0.50.4\lesssim\mu_{o}<$ 0.5. Therefore, while the critical phase is unstable in static Majorana chains and is thus unique to monitored Majorana chains, the bulk-edge correspondence is obscured since it corresponds to the gapless phase. This is reminiscent of the fact that the topological phase is usually well defined in gapped phases in static systems.
@@
Remarkably, however, we find that the topological invariant dynamically characterizes the different phases, including the critical gapless phase. Specifically, we consider the sample-averaged invariant $\chi_{T}$ for the timescale $T=\mathcal{O}(L)$. Figure 3(d) shows $\chi_{T=L}$ for different system sizes, from which we find three crossings with $\mu_{o}\simeq 0.4,0.5,$ and $0.6$. The locations of the first and last crossings coincide well with the transitions between the area-law gapped phases and critical gapless phase. Notably, this characterization is attributed to the divergent relaxation time of the invariant in the gapless phase. If we denote the timescale as $\tau_{\mathrm{relax}}$, it is at least longer than $\sim 1/z_{1}$, which increases faster than $L$ inside the gapless phase. Therefore, $\chi_{T=L}$ will not converge and almost stay zero for the gapless phase. This is contrasted to the gapped phases with $z_{1}=\mathcal{O}(L^{0})$, where $\chi_{T}$ rapidly converges to ±$\pm 1$ for $T=\mathcal{O}(L)$. Note that the crossing point at $\mu_{o}\simeq 0.5$ is different from the boundaries of the three phases. The above argument indicates that this crossing is a finite-size artifact, and $\chi_{T=L}$ will flatly become zero in the entire critical phase for $L\rightarrow\infty$.
Summary and outlook–
@@
We have shown that monitored Majorana circuits in the gapped phases exhibit the bulk-edge correspondence by investigating the Lyapunov spectrum, gapless edge modes, and topological invariants based on the fermion parity. To define the topological invariant, we have used a unique methodology where the trajectory under the APBC is defined with respect to that under the PBC, according to twisted measurement outcomes at the boundary. We have also shown that the critical phase corresponds to the gapless phase in terms of the Lyapunov spectrum and is dynamically characterized by the topological invariant in a timescale $t=\mathcal{O}(L)$.
@@
Our study opens the way to explore topological features in monitored systems through the Lyapunov analysis. An important future direction is to examine many-body systems, where our topological invariant based on the twisted measurement outcomes readily applies. For example, it is intriguing to analyze monitored circuits with $Z_{2}\times Z_{2}$ symmetry, for which measurements can stabilize an area-law phase with a symmetry-protected topological order of the cluster state [26, 38, 32]. Such a topological area-law phase will be characterized by a (nearly) fourfold degenerate Lyapunov spectrum under the OBC from gapless edge modes and a topological invariant measuring the change of one $Z_{2}$ charge by twisting boundary measurement outcomes for another $Z_{2}$ symmetry. Another promising direction is to extend our topological invariant to measurement-induced topological phases protected by $U(1)$ symmetry.
@@
Note added.— While this work was in its final stage, we learned about a recent paper that investigates a similar issue [90].